RC meeting and Run Plan, Friday November 4th
From SBS wiki
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Current Status
- 4 Run Plan After Configuration Change
- 5 Notices in Effect for Shift Crew
- 6 Shifts and Long Term Issues
- 7 Planned Accesses to the Hall
- 8 Special Requests
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
- Hall in restricted access for continued configuration change operations
- Beamline removed
- SBS move underway
- Possible updates from Jessie
- A final RF module needs replaced in BPM VME crate. Possible update from Simona.
- Compass field measurements were not successful. Gradient measurements will be used instead for kin 2. Possible updates from Bogdan.
- Oven heater variac controllers need replaced and tested. Possible updates from Todd.
- UVA optics assembly removed from optics enclosure and tested at EEL. Possible updates from Gordon.
- Helmholtz coil field measurements need to be conducted prior to moving to kin 3 configuration. Possible updates from Bill.
- Current Polarization: N/A (Target not installed at present)
- Hourly screenshots suspended until target installation. Lasers locked out.
- Any Updates?
- Hall in restricted access for remainder of configuration change
- Hall access is open to subsystem experts after 1700 and before 0800
- Please check with RC for access beyond DAQ bunker during the day
Shifts and Meetings
- Shift leader and target operator shifts are on stand-by for the weekend. At this stage, we will require SL and TO to return Monday Owl, but SL will likely be put on stand-by through Monday.
- Shall we conduct RC meetings over the weekend?
Experimental progress (for MCC planning meeting)
- 6.1% of accumulated charge for PRODUCTION data (13.55 Coulombs out of 232)
- 28% Experimental run time (40 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- Luminosity scans with GEMs
- Opportunistic BCM calibration (higher current) before pass change?
- INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
- Work and testing on LV supplies during opportunistic access.
- Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29
Run Plan After Configuration Change
May need updated 'standard' run plan for 3-pass data taking.
- 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour.
- 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization.
- 3.) Return to step 1.)
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: , If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
- Cross check the rate with what is expected: 
- Cross check with the expected ion chamber readings: 
Notices in Effect for Shift Crew
Be sure to follow these procedures:
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Shifts and Long Term Issues
- Shift Openings: Wednesday OWL, Thursday OWL. 
Planned Accesses to the Hall
In Restricted Access. Configuration change ongoing.
Any special requests?